The length of a rectangular field is forty meters more than the width. The area of the field is six thousand square meters. Write down the length in terms of the width and determine the length and the width of the rectangle.
Answer by addingup(3677)

W*L = 6,000
L = W+40
W(W+40)= 6,000
W^2+40W = 6,000
W^2+40W-6,000 = 0
(W-60)(W+100)
W = 60 or W = -100
Since we don't need a negative number, discard the 100 and try the 60:
W(W+40) = 6000
60(60+40) = 6000 Correct.